can you help with this quadratic equation
4x2 - 16 = 0

Answer by dolly(163)

4x^2 - 16 = 0
==> 4x^2 = 16
==> x^2 = 16/4
==> x^2 = 4
==> x = sqrt(4)
==> x = 2, -2